Perpetual Inventory Using LIFO Beginning inventory, purchases, and sales for Item PK95 are as follows: January 1 Inventory 96 units @ $32 5 Sale 77 units 11 Purchase 107 units @ $36 21 Sale 90 units Assuming a perpetual inventory system and using the last-in, first-out (LIFO) method, determine (a) the cost of merchandise sold on January 21 and (b) the inventory on January 31. a. Cost of merchandise sold on January 21 $ b. Inventory on January 31 $
Answer #1
Unit Unit Cost Total Cost
Jan 1 96 32 3072
Jan 5 77
Jan 11 107 36 3852
Jan 21 90

a) Cost of merchandise sold on January 21 = 90*36 = 3240

b) Ending inventory = (17*36+19*32) = 1220